# ChartCoder: Advancing Multimodal Large Language Model for Chart-to-Code Generation

### v1

@@ 5.4 Analysis
We further evaluate the role of code in the chart understanding task. We use MiniCPM-Llama3-V2.5 to evaluate two input forms, Image only and Image with Code, on the MMC True/False benchmark Liu et al. (2023a). The result in Table 6 shows that using code helps the model better understand chart details, especially the chart types and the data they contain. A case study is shown in Figure 6.
Benchmark Image Image+Code $\triangle$
MMC T/F 0.76 0.81 6.5%
Table 6: Comparison of the impact of using ChartCoder-generated code as auxiliary information on the MMC True/False benchmark.
5.5 Conclusion
@@
Our study is comprehensive but has certain limitations that we aim to address in future research. Due to constraints in computational resources, we only trained ChartCoder with 7B parameters, which has demonstrated sufficiently good results for now. A larger model could potentially achieve even better performance. Future work may focus on exploring more complex and diverse charts and codes while also experimenting with other image types, such as HTML, to develop a comprehensive multi-modal code large language model.
@@ A.3 More Ablation Studies
A.4 Chart2Code-160k Analysis
We count the proportion of different charts in the Chart2Code-160k dataset in Table 8.

### v2

@@ 5.4 Analysis
We further evaluate the role of code in the chart understanding task. We use two MLLMs to evaluate two input forms, Image only and Image with Code, on the MMC True/False benchmark Liu et al. (2023a). The result in Table 7 shows that using code helps the model better understand chart details, especially the chart types and the data they contain. A case study is shown in Figure 5. Also, we utilize LLM to evaluate the readability of ChartCoder output code, and details are in the Section A.4.
6 Conclusion
@@
Our study is comprehensive but has certain limitations we aim to address in future research. Due to constraints in computational resources, we only trained ChartCoder with 7B parameters, which has demonstrated sufficiently good results for now. A larger model could potentially achieve even better performance. Future work may explore more complex and diverse charts and codes while experimenting with other image types like HTML to develop a comprehensive multi-modal code large language model.
@@
Acknowledgment
The work is initiated and supported by the AI9Stars Team.
@@ A.3 More Ablation Studies
A.4 Output Code Analysis
To evaluate the output code readability, we conduct an ablation experiment, utilizing gpt-4o-2024-08-06 to evaluate the output code readability. We evaluate four aspects of the generated code, including Naming Conventions, Code Structure, Comments, and Logical Clarity, with a total score of 100. We choose the generated code from the ChartMimic task (ChartCoder output) and the ground truth code (human-annotated) in the ChartMimic dataset. The results are as shown in Table 9. We also evaluate the error types on ChartMimic direct generation tasks with code and general LLMs as the language backbone. The results are shown in Figure 4.
Dataset Source Code Readability
Mean $\mu$ SD $\sigma$
Qwen2-VL-7B Output Model generated 82.48 6.81
ChartCoder Output Model generated 85.22 6.78
ChartMimic Source Human written 87.66 4.30
Table 9: Performance Comparison of model outputs and human-written sources. SD is the abbreviation ddfor standard deviation.
@@ A.4 Output Code Analysis
A.5 Chart2Code-160k Analysis
We count the proportion of different charts in the Chart2Code-160k dataset in Table 10. Also, we utilize gpt-4o-2024-08-06 to evaluate the quality of the charts in the Chart2Code-160k and compare them with the real-world chart. The prompts are shown in Figure 8.
@@ A.5 Chart2Code-160k Analysis
Prompt for Code Readability Evaluation Please score the code’s readability based on the following four aspects. Each aspect is worth 25 points, for a total of 100 points. Naming Conventions (25 points) Score: [X]/25 Explanation: [Provide a brief explanation of how well the variable, function, and class names convey their purpose and whether the naming style is consistent across the codebase.] Code Structure (25 points) Score: [X]/25 Explanation: [Explain whether functions are concise, whether the code uses indentation and blank lines appropriately, and whether the code is modularized effectively.] Comments (25 points) Score: [X]/25 Explanation: [Discuss the clarity and appropriateness of the comments, and whether functions/methods have proper documentation comments explaining inputs, outputs, and functionality.] Logical Clarity (25 points) Score: [X]/25 Explanation: [Evaluate the intuitiveness of the code, whether it’s easy to understand, and whether the control flow is simple and avoids unnecessary complexity.] Total Score: [X]/100 Summary: [Provide a brief overall assessment of the code’s readability, pointing out strengths and potential areas for improvement.]
Figure 7: Prompt for dataset quality evaluation.
Prompt for Chart Quality Evaluation You are a professional chart analyser. Please evaluate the image based on the following four criteria: aesthetics, readability, reproducibility, and data presentation simplicity. Provide a score for each criterion and include an overall score along with a brief evaluation. Scoring Criteria and Requirements: Aesthetics (25 points) Requirements: The chart design should be simple and clear, avoiding complex decorations, and should effectively communicate information. Colors should be harmonious and have high contrast, making it easy to differentiate between different data groups. Legends and labels should be clear, with appropriately sized fonts, avoiding visual clutter. Scoring: [X]/25 Readability (30 points) Requirements: The chart should have clear titles, axis labels, and legends, enabling quick communication of the main message. Data curves or point annotations should avoid being overly dense or overlapping, maintaining good readability. The overall layout should follow a logical structure without any confusing elements. Scoring: [X]/30 Reproducibility (30 points) Requirements: The chart design should be easy to replicate using common tools. Data availability is critical: even if the design is simple, missing context or data should result in point deductions. Data should be provided with clear sampling methods, units, and formats, enabling others to recreate the chart from scratch. The presentation of data should align logically with the chart design, avoiding overly customized or complex elements. Scoring: [X]/30 Data Presentation Simplicity (15 points) Requirements: Data presentation should be concise, avoiding redundant information. Data should be displayed in an intuitive way, without excessive curves, points, or annotations. High-scoring charts should focus on the main data being presented, avoiding decorative or unrelated information. Scoring: [X]/15
Figure 8: Prompt for dataset quality evaluation.